OpenStudy (anonymous):

help to prove sinAsin(60-A)sin(60+A) = 1/4sin3A

OpenStudy (raden):

use some identity trigonometry formula here, like : sin(A+B) = ... (cosA)^2 = ... sinAsinB = .... etc.

OpenStudy (raden):

for sin(3A) = sin (A+2A) = ... can u expand that ?

OpenStudy (anonymous):

sinAcos2A+cosAsin2A

OpenStudy (raden):

ok, right.. from that convert sin2A = 2sinAcosA, what happend next ?

OpenStudy (anonymous):

sinAcos2A+cosA(2sinAcosA) sinA [cos2A+2cos^2A] sinA[cos2A+2(1-sin^2A)] sinA[cos2A+2-2sin^2A] sinA[ 1-2sin^2A +2-2sin^2A] sinA [3-4 sin^2A]

OpenStudy (raden):

OK, correct.. so, to right side we have (1/4)sin3A = (1/4)sinA [3-4 sin^2A], right ?

OpenStudy (anonymous):

I get the way!!! sin A sin (60-A)sin(60+A) = sinA[-1/2(cos120-cos2A)] = sinA{-1/2[-1/2-(1-2sin^2A)]} = sin A[3/4-sin^2 A] = 1/4 sin A [ 3-4sin^2 A] = 1/4sinA[ 1-2sin^2A +2-2sin^2A] = 1/4sinA[cos2A+2-2sin^2A] = 1/4sinA[cos2A+2(1-sin^2A)] = 1/4sinA [cos2A+2cos^2A] = 1/4sinAcos2A+cosA(2sinAcosA) = 1/4sinAcos2A+cosAsin2A = 1/4sin (A+2A) = 1/4sin3A (proven) THANKS SO MUCHHH!!!

OpenStudy (raden):

woww, congrate.. my pleasure ^^